### v2

# Imagination Augmented Generation: Learning to Imagine Richer Context for Question Answering over Large Language Models

Abstract
Retrieval-Augmented-Generation and Gener-ation-Augmented-Generation have been proposed to enhance the knowledge required for question answering over Large Language Models (LLMs). However, the former depends on external resources, and both require incorporating the explicit documents into the context, which results in longer contexts that lead to more resource consumption. Recent works indicate that LLMs have modeled rich knowledge, albeit not effectively triggered or activated. Inspired by this, we propose a novel knowledge-augmented framework, Imagination-Augmented-Generation (IAG), which simulates the human capacity to compensate for knowledge deficits while answering questions solely through imagination, without relying on external resources. Guided by IAG, we propose an imagine richer context method for question answering (IMcQA), which obtains richer context through the following two modules: explicit imagination by generating a short dummy document with long context compress and implicit imagination with HyperNetwork for generating adapter weights. Experimental results on three datasets demonstrate that IMcQA exhibits significant advantages in both open-domain and closed-book settings, as well as in both in-distribution performance and out-of-distribution generalizations 111 Our code will be available at https://github.com/Xnhyacinth/IAG.
1 Introduction
Knowledge-intensive tasks like question answering (QA) necessitate access to extensive world and domain knowledge Berant et al. (2013); Joshi et al. (2017); Kwiatkowski et al. (2019). Recently, Large Language Models (LLMs) have displayed notable competencies in almost every task and industry within the “pre-train, prompt, and predict” paradigm Liu et al. (2023b). However, LLMs lack the sufficient capability to independently handle knowledge-intensive tasks Yu et al. (2023) and usually generate hallucinations Zhao et al. (2023).
Figure 1: Compared with RAG (Top) and GAG (Middle), the proposed IAG (Bottom) eschews external resources, utilizing solely LLMs to imagine a shorter explicit document and more flexible implicit adapters.
In recent years, to alleviate the issue of hallucinations in LLMs and improve performance in knowledge-intensive tasks such as QA, researchers have proposed numerous knowledge-augmented methods for LLMs, which mainly include two categories: Retrieval-Augmented-Generation (RAG) Guu et al. (2020) and Generation-Augmented-Generation (GAG) Abdallah and Jatowt (2023). RAG (Top part of Figure 1) retrieves related documents from external resources (e.g., auxiliary tools and domain documents) and then sends those retrieved documents and the question together into LLMs Izacard and Grave (2021). RAG has demonstrated formidable performance on several tasks Lewis et al. (2020). However, RAG not only requires presupposed external resources but also necessitates more computational resources and longer processing times Xu et al. (2023a). Take typical RAG method FiD Izacard and Grave (2021) as an example, the required computing resources and inference time continue to increase as the number of retrieved documents increases. While retrieving 100 documents requires processing over 12k tokens, it will result in an exceeding 100×100$\times$100 × prompt length (decades times of GPU consumption) and an over $100^{2}\times$ computational time Liu et al. (2023a).
To avoid dependence on external resources, GAG Yu et al. (2023) has been proposed and can utilize LLMs like InstructGPT Ouyang et al. (2022) to generate the relevant documents (Middle part of Figure 1). However, it requires additional financial costs (e.g., API calls) and still demands substantial computational resources and time. Additionally, both RAG and GAG utilize more explicit external resources (symbolized documents), and the quality of the obtained content significantly impacts downstream tasks Li et al. (2023); Shaier et al. (2024). For example, Gao et al. (2024) indicates that noise in the documents negatively affects the performance. Therefore, there is an urgent need to explore new knowledge-augmented methods.
In fact, LLMs inherently contain rich knowledge and possess significant potential for resolving knowledge-intensive tasks Bhagavatula et al. (2020). Enhancing the performance of specific tasks can be achieved by better activating relevant knowledge or expanding memory capacity without relying on external resources. For example, simply repeating the question twice Xu et al. (2023b), just reviewing and consolidating knowledge by appending a straightforward prompt “As far as I know” Yao et al. (2023), and using visual-language models to imagine images Tang et al. (2023), all approaches enhance the performance of LLMs on downstream tasks. Inspired by this, we introduce a novel knowledge-augmented framework Imagination-Augmented-Generation (IAG) for LLMs, which simulates the human capacity to compensate for knowledge deficits solely through imagination in QA. As shown in the bottom part of Figure 1, for resolving knowledge-intensive tasks, IAG utilizes solely LLMs to imagine a shorter explicit document and more flexible implicit adapters.
Within the framework of IAG, we introduce an imagine richer context method for question answering (IMcQA). To sufficiently utilize the inherent knowledge of LLMs, we design two main modules to activate the various potential knowledge modeling in LLMs and obtain a richer context. Specifically, the explicit imagination module first uses symbol distillation to obtain the compressed context and then guides LLMs in generating a short and useful dummy document. Subsequently, the implicit imagination module utilizes the proposed HyperNetwork to generate LoRA weights to activate the task-processing ability of LLMs. Unlike the LoRA Hu et al. (2021) stores task knowledge and ability in modules, the HyperNetwork learns to imagine hidden knowledge for each question.
We evaluate the proposed IMcQA to various LLMs, including T5 Roberts et al. (2020a) and Llama2 Touvron et al. (2023). The experimental results across three QA datasets indicate that the proposed method yields performance gains while reducing computational expenses and time. Notably, it even outperforms baseline methods that retrieve and generate knowledge under the same document settings. In conclusion, the contributions of this paper are summarized as follows:
  • We propose a new knowledge augmentation framework IAG to fully leverage the LLMs’ intrinsic knowledge more efficiently without relying on external resources.
  • We propose a novel QA method IMcQA that employs two main modules (explicit imagination and implicit imagination) to better utilize the knowledge stored in the LLMs and obtain richer context in QA.
  • Experimental results indicate that the proposed method successfully activates the relevant internal knowledge of LLMs. IMcQA exhibits significant advantages in both open-domain and closed-book settings, as well as in both in-distribution performance and out-of-distribution generalizations.
2 Related Work
This paper mainly utilizes context compression, hypernetworks and knowledge distillation to achieve knowledge enhancement. The following will elucidate pertinent research across four facets.
Knowledge Enhancement has usually been adopted to alleviate the issue of insufficient knowledge in LLMs. There are two main methods: RAG Sun et al. (2019) and GAG Abdallah and Jatowt (2023). The typical RAG method FiD Izacard and Grave (2021) retrieves the documents from an external knowledge base to answer questions. As LLMs can be considered a knowledge base, several studies Liu et al. (2022) propose to extract knowledge from LLMs (e.g., GPT-3). For example, Yu et al. (2023) generates 10 relevant documents for world knowledge according to the question. However, RAG needs the related external resources, and both RAG and GAG still need to obtain and utilize verbose explicit long contexts. Recently, there have been some methods to enhance the LLMs’ ability through simulating human imagination of visual information. But they use existing visual-language models Tang et al. (2023); Akter et al. (2024), while we prefer self-imagination to augment knowledge. Besides, they have not fully leveraged the parameterized knowledge within the models Xu et al. (2023b); Kazemnejad et al. (2023). In this paper, the proposed method for augmenting knowledge not only obviates the need for external resources but also enhances the efficiency of extracting and activating internal knowledge within LLMs.
Context Compression has often been used to improve the efficiency of LLMs in processing long contexts. Recent studies Mu et al. (2023) propose that long contexts be condensed into summary vectors (soft prompts) to ensure their effective utilization by LLMs. Simultaneously, some studies Jiang et al. (2023) believe that information redundancy in lengthy texts and information entropy can be utilized to compress the contexts Li et al. (2023). Unlike them, this paper is devoted to awakening the long-context modeling ability of LLMs. By learning an Imagine Model that can generate compressed contexts, the QA model that operates on short contexts can also possess a rich contextual understanding akin to the QA model designed for processing longer contexts.
Knowledge Distillation is a technique where a smaller model learns to mimic the predictions of a larger model, aiming to retain performance while reducing computational resources Hinton et al. (2015). Recent studies West et al. (2022) present symbolic knowledge distillation, a process that facilitates knowledge transfer from a teacher model via extracting training data to subsequently train a student model. In this paper, the process of obtaining compressed context during the explicit imagination resembles a form of symbolic distillation. Regarding training, our emphasis lies in distilling the long-context modeling abilities of LLMs.
Hypernetworks is designed to reduce the number of parameters Ha et al. (2016), i.e., a small neural network generates parameters for another big neural network. Recent studies Phang et al. (2022); Ivison et al. (2023) have explored the enhancement of model performance in zero- and few-shot settings through meta-learning involving hypernetworks. We utilize hypernetworks to acquire implicit imagine capabilities by dynamically generating LoRA for efficiency and generalization.
3 Method
In this section, we introduce the detailed method of IMcQA to activate LLMs’ intrinsic knowledge and obtain a richer context for QA. The fundamental premise underlying this method is that QA with a richer context yields greater performance. Consequently, diverse methods are employed for questions lacking in richer contexts to activate knowledge within LLMs to replicate comparable effects to those achieved with richer contexts.
Specifically, IMcQA comprises two main modules. Explicit imagination with long context compression learns to imagine a short dummy document (§ 3.2). And implicit imagination with the HyperNetwork models hidden knowledge that learns a shared knowledge feature projection across questions (§ 3.3). The HyperNetwork is trained to generate lightweight LoRA modules, aiming to align the question and the internal knowledge. Besides, there is long context distillation in training, which learns the teacher’s rich representations to compensate for missing knowledge in imagination (§ 3.4).
3.1 Formulation
The formulation of our task follows RAG for QA Guu et al. (2020). Let $\mathcal{V}^{*}$ denote the infinite set, encompassing all potential strings over the tokens in vocabulary $\mathcal{V}$, and this includes the empty string. An instance within a QA dataset is defined as a triplet $(q,a,c)$ comprising question $q$, answer $a$, and context $c$, where $q,a,c\in\mathcal{V}^{*}$. Conventionally, the context $c$ is drawn from the knowledge corpus $\mathcal{Z}$, like Wikipedia, whereby $\mathcal{Z}\subset\mathcal{V}^{*}$.
The goal of QA is to learn a distribution function $p(a|q)$. In a closed-book setting, LLMs directly encode the given question $q$ and generate the answer $a$ Roberts et al. (2020b). However, employing a direct approach of requesting models to output answers frequently results in poor performance, primarily attributable to the omission of a substantial amount of world knowledge. Therefore, a popular approach is the open domain setting, which marginalizes $p(a|q,c)$ over contexts $c$. Additional background details are available in A.1.
Figure 2: Overview of IMcQA method. In the inference phase, for each question, the Imagine Model imagines a short dummy document based on the question and the HyperNetwork generates specific LoRA weights. During training, there are two stages: the first stage is the pre-training of the Imagine Model, aiming at its ability to imagine a short dummy document based on the question, and the second stage is the HyperNetwork fine-tuning using long context distillation (§ 3.4) to learn a map from the question to the LoRA weights.
3.2 Explicit Imagination with Compress
To get the context $c$, we utilize LLMs to imagine a short dummy document, which can mitigate knowledge corpus error Lee et al. (2023) by considering potentially useful contexts. In the view of compression, we greatly reduce input length, minimize noise, and elevate the salience of essential tokens.
As shown in the left part of Figure 2, to help LLMs fully utilize the knowledge and imagine compressed text, we first pretrain it on our collected question-compressed document pairs. By leveraging symbolic distillation, we employ the long-context compression method LongLLMLingua Jiang et al. (2023) to condense a large corpus of retrieved documents. These compressed texts $c^{\prime}$ then serve as fine-tuning data alongside specific prompts $p_{q}$ (A.2) and question-answer pairs for the Imagine Model $G_{\theta}$ ($\theta$ represents the model’s parameter), which guides the model to think about its knowledge and imagine a short dummy document:
$d=G_{\theta}(p_{q}(q;c^{\prime}))$ (1)
where $d$ is the dummy document generated from the Imagine Model. This process enables LLMs to conceive compressed knowledge that robustly parallels the question’s knowledge requirements.
3.3 Implicit Imagination with HyperNetwork
We advance upon LoRA Hu et al. (2021) by suggesting the implementation of the HyperNetwork, which does not directly optimize the LoRA module but generates specific LoRA adapter weights using the inputs for QA (bottom part of Figure 2). This is akin to repeating the question in the prompt Xu et al. (2023b) and incorporating certain topical cues to stimulate the model’s recall of relevant questions Wang et al. (2023). However, the distinction lies in the fact that they serve as wake-up features, whereas we are generating model parameters.
Figure 3: The Architecture of HyperNetwork. HyperNetwork generates LoRA adapter weights for each question. During training, only HyperNetwork, FFN, and Norm weights are updated.
The HyperNetwork architecture for generating LoRA weights is exhaustively outlined in Figure 3. Specifically, $D_{k}^{q}$ and $U_{k}^{q}$ represent the low-rank projections of layer $k$ correlated with the $\mathcal{Q}$, while $D_{k}^{v}$ and $U_{k}^{v}$ correspond to those associated with the $\mathcal{V}$. HyperNetwork represented as $g_{D}$ and $g_{U}$, takes $\textit{concat}(f,i_{k}^{\{q,v\}})$ as input, where $f$ is the feature vectors that use the model’s encoder to obtain and using whitening algorithm Su (2021) for dimensionality reduction, $i_{k}^{\{q,v\}}\in\{0,\ldots,2\times\#\text{blocks}\}$ signifies the positional embedding differentiating between layers and between $\mathcal{QV}$. Each HyperNetwork is defined by weights $W_{d}$ and $W_{u}$ which represent the down and up projections respectively. Finally, the HyperNetwork equations for $D^{\{q,v\}}$ and $U^{\{q,v\}}$ can be expressed as:
$f=\text{whitening}(\text{Encoder}(q;d))$ (2)
$x=\text{concat}(f,i_{k}^{\{q,v\}}~{}|~{}i_{k}^{q}=2k,i_{k}^{v}=2k+1)$ (3)
$D^{\{q,v\}},U^{\{q,v\}}=g_{D}(x),g_{U}(x)$ (4)
where the Encoder represents the encoder of the model, whitening is a dimensionality reduction algorithm, and concat means to splice the content. $g_{D}$ and $g_{U}$ denote the descending and ascending dimensions of HyperNetwork. More formally,
$g(x)=\text{MM}(\text{ReLU}(\text{MM},W_{d}),W_{u})$ (5)
where MM stands for matrix multiplication, ReLU is a activation function.
3.4 Training with Long Context Distillation
Within the framework of knowledge distillation, components such as hidden representations Jiao et al. (2020), attention dependencies Wang et al. (2020), and relations among representations Park et al. (2021) are regarded as valuable knowledge for transfer. In this paper, we consider long context distillation (LCD) as the contextualized knowledge that mainly guides the student.
Specifically, the teacher model FiD Izacard and Grave (2021), which utilizes longer contextual inputs and theoretically contains more information (richer context). It will activate more specific internal knowledge and serve as a supervisory model. The teacher model assists the student model T5 Roberts et al. (2020a), which has the same size as the teacher and leverages short contextual inputs. This aids in activating richer feature representations and knowledge. The optimization objective for the student model at each mini-batch $z_{r}=(x_{r},y_{r})$ is:
$\begin{split}\mathcal{L}_{\text{s}}(\theta_{s},\theta_{t},z_{r})=\alpha\mathcal{L}_{\text{ce}}(y_{r},S(x_{r};\theta_{s}))\\ +(1-\alpha)\mathcal{L}_{\text{ce}}(T(x_{r};\theta_{t}),S(x_{r};\theta_{s}))\end{split}$ (6)
where we have a teacher model denoted as $T(\cdot;\theta_{t})$ and a student model denoted as $S(\cdot;\theta_{s})$. The corresponding model parameters are $\theta_{t}$ and $\theta_{s}$.
As shown in the right of Figure 2, we perform an additional representation alignment for better knowledge transfer. In our distillation, both teacher model and student model have the $L$ layers, we feed the text into them and can obtain the corresponding output hidden states $\{H_{l}^{t}\}_{l=0}^{L}$, $\{H_{l}^{s}\}_{l=0}^{L}$, and attention matrices $\{A_{l}^{t}\}_{l=1}^{L}$, $\{A_{l}^{s}\}_{l=1}^{L}$. We suppose the student’s $l$-th layer is aligned with the teacher’s $l$-th layer, then the outputs of the student (i.e., $H_{l}^{s}$ and $A_{l}^{s}$) should be close to the teacher’s (i.e., $H_{l}^{t}$ and $A_{l}^{t}$). For aligning hidden states, following Park et al. (2021), we use cosine distance COS to calculate the proximity between the hidden states of the teacher and the student:
$\mathcal{L}_{\text{hid}}=-~{}\text{COS}(H_{l}^{s},H_{l}^{t})$ (7)
While for aligning attention dependencies, we follow Jiao et al. (2020) to optimize the mean square error (MSE) between the attention matrices of the teacher and the student:
$\mathcal{L}_{\text{attn}}=-~{}\text{MSE}(A_{l}^{s},A_{l}^{t})$ (8)
The overall objective for knowledge transfer is:
$\mathcal{L}_{\text{align}}(H_{l}^{s},H_{l}^{t},A_{l}^{s},A_{l}^{t})=\mathcal{L}_{\text{attn}}+\mathcal{L}_{\text{hid}}$ (9)
The overall objective for training IMcQA is the weighted sum of the two objectives:
$\mathcal{L}=\mathcal{L}_{\text{s}}+\lambda\mathcal{L}_{\text{align}}$ (10)
4 Experiment
In this section, we conduct experiments to demonstrate the effectiveness and efficiency of IAG and IMcQA on QA. The experiment mainly answers four research questions (RQs):
RQ1: Can IAG achieve knowledge augmentation for answering questions over LLMs? (§ 4.4)
RQ2: Does our method have good knowledge activation and generalization abilities? (§ 4.5)
RQ3: Does IAG have advantages in effectiveness and efficiency compared to RAG and GAG? (§ 4.7)
RQ4: What is the role of explicit and implicit imagination modules in IAG and IMcQA? (§ 4.8)
Models Reader Params # Documents NQ TriviaQA WebQ
# closed-book setting
T5-l Roberts et al. (2020a) 770M 0 28.5* 28.7* 30.6*
T5-xl 3b 0 28.30 33.92 34.43
IMcQA-l (Ours) 770M 0 29.32 30.11 32.68
IMcQA-xl (Ours) 3b 0 29.59 35.71 37.40
# Retrieval Augmented Generation
DPR Karpukhin et al. (2020) 110M 100 41.5* 56.8* 41.1*
RAG Lewis et al. (2020) 400M 10 44.5* 56.1* 45.2*
FiD-l Izacard and Grave (2021) 770M 10 46.7* 61.9* 48.1*
FiD-xl 3b 10 50.1* 66.3* 50.8*
# Generation Augmented Generation
GENREAD-l Yu et al. (2023) 770M 10$\dagger$ 40.3* 67.8* 51.5*
GENREAD-xl 3b 10$\dagger$ 42.6* 69.6* 52.6*
# Imagination Augmented Generation (Ours)
IMcQA-l 770M 1$\dagger$ 42.32 65.48 45.28
IMcQA-xl 3b 1$\dagger$ 46.51 68.38 50.45
IMcQA-l 770M 10 49.92 69.67 51.52
IMcQA-xl 3b 5$\ddagger$ 50.87 70.34 52.78
Table 1: QA performances (%) of different methods on three datasets. The first part (closed-book setting) indicates that only utilize questions; The latter three parts utilize explicit documents. The best results are in bold. * means that those results are from existing papers, $\dagger$ denotes that the documents were generated ($\ddagger$ indicates that the number of documents is reduced due to insufficient memory for distillation). More results can be seen in Appendix A.5.1.
4.1 Datasets
We evaluate the proposed approach on three public question answering datasets: NaturalQuestions (NQ) Kwiatkowski et al. (2019), WebQuestions (WQ) Berant et al. (2013) and TriviaQA (TQA) Joshi et al. (2017). To evaluate the model performance, we use the exact match (EM) score for evaluating predicted answers Rajpurkar et al. (2016). We provide dataset details in the Appendix A.4.
4.2 Baselines
[…]
For the zero-shot settings of LLMs ($\geq$ 3B), we use Llama2-7B and Llama2-13B Touvron et al. (2023) as the basic model. We evaluate with four diverse settings: without retrieval, with retrieval, with LoRA, and using the proposed IMcQA.
4.3 Implementations
In the pretraining stage, the Imagine Model initialized with T5-large utilizes the generated question-compressed pairs. During the second stage, the teacher model employs a FiD reader with different sizes (FiD-l and FiD-xl) that are fine-tuned on the training split of target datasets. The student model freezes the backbone and updates solely the HyperNetwork, the feedforward neural network (FFN), and the normalization layers Chen et al. (2023). More implementation details and experimental findings are in the Appendix A.3.
4.4 Main Results
Table 1 shows the performance results, full results including T5-Base are in the Appednidx A.5.1.
As shown in Table 1, when juxtaposed with closed-book models, RAG, and GAG methods, our proposed IAG framework IMcQA method exhibits state-of-the-art performance with the equivalent magnitude of document count and model size.
In the closed-book setting (in the upper part of the table), our method outperforms the baseline by an average of +2% EM score, indicating its excellence in utilizing internal knowledge with imagination. It’s especially noteworthy that as the model size expands, the performance advantages of the imagination become ever more evident.
The following three parts of Table 1 show the experimental results under the open domain setting. Although our method only deals with one short dummy document, it can still achieve results similar to or better than the RAG and GAG methods, which handle 10 documents. The findings reveal that IMcQA exploits imagined condensed text to strike a balance between efficiency and overhead. Moreover, when IMcQA utilizes 10 retrieved documents, it supersedes the performance of FiD and GENREAD with the EM average of +2.26% and +3.06%.
4.5 Out-Of-Distribution (OOD) Performance
To further demonstrate the generalizations of the IMcQA method and the importance of HyperNetwork, we also evaluate its performance in out-of-distribution (OOD) generalizations. Table 2 shows the IID and OOD performance of FiD, and IMcQA methods with different document settings when training on NQ (From NQ generalization to the other two datasets). Full OOD results of three datasets are shown in the Appendix A.5.2.
It is patently clear that an increment in document provision leads to better OOD performance, likely due to the presence of answer-oriented content within these documents. Remarkably, IMcQA can come within a relatively narrow 5% gap of FiD, even when utilizing a single imagined document as opposed to 10 retrieved ones.
Simultaneously, IMcQA generally showcases superior performance in OOD when provided with 10 retrieved documents. This superiority can be traced back to the pivotal role played by HyperNetwork in generating LoRA adapters’ weights based on questions. This equips models with the capability to invoke and access internal knowledge based on context-specific discourse rather than confining to resolving distinct questions.
4.6 Zero-shot Results on LLMs
Figure 4 and Figure 5 illustrate the zero-shot results for LLMs implementing IMcQA. This research seeks to explore the possibility of enhancing LLMs via IAG. Due to the high computational demands of training, we only fine-tuned the HyperNetwork on a mixed dataset without LCD in this experiment and evaluated performance in a zero-shot setting. Detailed prompt information can be found in the Appendix A.2.
Models # Docu- NQ
ments NQ TQA WQ
T5 0 22.16 3.18 4.12
IMcQA 0 23.89 6.21 10.94
IMcQA 1$\dagger$ 40.14 46.61 18.92
FiD 10 46.81 53.93 24.02
IMcQA 10 47.01 55.74 24.13
T5-l 0 28.5* 3.18 4.12
IMcQA-l 0 29.32 10.17 14.06
IMcQA-l 1$\dagger$ 42.32 54.80 22.05
FiD-l 10 46.7* 57.93 25.12
IMcQA-l 10 49.92 60.03 25.79
Table 2: OOD results. The primary row in the table header delineates the dataset trained, while the underscored secondary row demonstrates the in-distribution performance (IID).
We discerned that Llama2’s performance can be enhanced by imagining knowledge autonomously. While leveraging explicit imagined context could amplify the average EM +1%, this is not as significant as the improvement achieved by retrieving 10 documents, indicating the limitations of relying solely on prompt cues for triggering corresponding knowledge. IMcQA can enhance knowledge via two main imagination processes, escalating EM by +15.33% for NQ, +11.97% for TQA, and +16.38% for WQ. With IAG, Llama2-7B demonstrated an average improvement of +14% across the three datasets. This trend is also observed in Llama2-13B’s results. This implies that even in zero-shot settings, our method can still offer substantial benefits to LLMs. More results can be seen in A.5.3.
4.7 Training Cost and Inference Speed-up
Models Training Params # Documents # Avg Tokens Inference Time GPU Memory
T5 220M 0 19.8 79.8s 2828M
IMcQA 139.3M 0 19.8 82.3s 2710M
IMcQA 139.3M 1 522.1 214.6s 2882M
FiD 220M 10 1748.3 683.3s 4358M
GENREAD 220M 10 1912.5 704.8s 4412M
FiD 220M 100 16625.7 1293.2s 19048M
Table 3: Training and inference cost on the NQ. The backbone model is T5-Base.
We proceeded to measure the inference speed, documented in GPU time, and training time for 5000 steps on the NQ dataset, using T5-Base. The experiments were conducted on a single RTX 3090 GPU, maintaining a standard batch size of 8 during training and 1 during inference.
As evident from Table 3, the proposed method’s advantage lies in its diminished requirement for parameter updates, which can be attributed to the shared HyperNetwork’s utilization that generates LoRA adapters, thereby negating the necessity of individual LoRA adapters’ setup. Despite the lack of a training advantage owing to distillation constraints, our approach achieves efficient reasoning through an extremely lightweight design. Compared to the other two methods, the token processing count is significantly decreased, while either outperforming them or showing negligible differences in performance. This represents an optimal trade-off between efficiency and computational demand. Moreover, unlike GAG, our approach incurs no financial costs associated with API calls, and the reduced model size facilitates faster generation.
Figure 4: Zero-Shot results (EM, %) of Llama2-7B on three open-domain QA datasets.
4.8 Ablation Experiment
In this study, we introduced two key imagination processes to stimulate LLMs’ internal knowledge: Explicit Imagination (EI) and Implicit Imagination (II). We particularly examined the influence of different imagination types on performance.
Figure 6 demonstrates that both EI and II are important for IMcQA. Omitting either one results in a considerable reduction in performance, with a drop exceeding 10% observed when EI is neglected. This is harmonious with the initial observation that performance improvement becomes more noticeable when relevant documents are available, thus underscoring EI’s superiority.
Figure 5: Zero-Shot results (EM, %) of Llama2-13B on three open-domain QA datasets.
The outcomes of Long Context Distillation (LCD) and the application of EI in the HyperNetwork also make marginal contributions to the overall results. This validates the previous assertion that a more extensive context tends to optimize performance, although with limited gains.
Figure 6: Ablation experiment results (%) of T5-Base on three open-domain QA datasets.
5 Conclusion and Future Work
This study proposes a novel knowledge-augmented strategy for LLMs, namely Imagine Augmented Generation (IAG), and a specific method IMcQA for open domain question answering. The proposed method effectively activates and utilizes intrinsic knowledge within LLMs through two imaginations: explicit imagination, and implicit imagination. Experimental results demonstrate a significant improvement in QA performance while remaining relatively lightweight. Although the main focus of this method is on one specific task, we believe these findings can offer a novel perspective on how to better harness the potential of LLMs. In the future, we plan to apply IAG to more NLP tasks and explore multimodal knowledge-augmented generation.
Limitations
While this study has demonstrated significant achievements in QA tasks, there are notable limitations:
Tasks.   The proposed methods in the study are specialized specifically for QA. It remains unknown how effective they would be in other types of knowledge-intensive tasks, such as fact-checking or dialogue systems. Further validation is needed to assess the generalizations and applicability of this approach.
Multimodal.   We have only considered imagined text and hidden representations. In future work, it is imperative to explore multimodal information including the impact of imagining images on performance.
Method.   Our method relies on the knowledge learned by LLMs in the pre-training phase, which may limit the model’s ability to quickly adapt to new information. The dependency on internal knowledge activation in IAG may lead to a less transparent decision-making process in the model, making it challenging to explain the logic behind the generated answers. In the future, there is a need to continue exploring adaptive knowledge enhancement methods to optimize results further.
Ethical Considerations
In this paper, we proposed a novel knowledge enhancement method aimed at leveraging the knowledge of LLMs. However, LLMs may generate inappropriate or discriminatory knowledge. Our approach does not introduce ethical concerns. The datasets we used are public, and there are no privacy issues.
Appendix A Appendix
A.1 Background
Our task formulation follows retrieval augmented models for QA Guu et al. (2020); Sachan et al. (2021). Let $\mathcal{V}^{*}$ denote the infinite set, encompassing all potential strings over the tokens in vocabulary $\mathcal{V}$, and this includes the empty string. An instance within a QA dataset is defined as a triplet $(q,a,c)$ comprising question $q$, answer $a$, and context $c$, where $q,a,c\in\mathcal{V}^{*}$. Conventionally, the context $c$ is drawn from the knowledge corpus $\mathcal{Z}$, like Wikipedia, whereby $\mathcal{Z}\subset\mathcal{V}^{*}$.
The goal of QA is to learn a distribution function, represented as $p(a|q)$, wherein the models decode a string $a$ that serves as an abstractive answer to a given query $q$. In a closed-book setting, LLMs directly encode the given question and predict the answer Roberts et al. (2020b). Specifically, considering the context $c$ as the empty string, the reliance is solely on the model parameters, i.e., $\hat{a}=\arg\max_{a\in\mathcal{V}^{*}}p(a|q,\theta)$, where $\theta$ represents the LLMs’ parameters. However, employing a direct approach of requesting models to output answers frequently results in subpar performance, primarily attributable to the omission of a substantial amount of world knowledge during the process. Therefore, a popular approach is open domain setting, which marginalizes $p(a|q,c)$ over contexts $c$ in the knowledge corpus Lewis et al. (2020); Sachan et al. (2021) or generated from models Yu et al. (2023). Given the computational infeasibility of calculating probabilities for all contexts, $p(a|q,c)$ is approximated to the sum of probabilities for top $k$ contexts, i.e., $p(a|q,c)=\sum\limits^{c_{i}\in c}\limits_{c\in TopK(q)}p(a|q,c_{i})p(c_{i}|q)$, where $Topk(q)$ denotes the set of resulting top $k$ passages after the retrieval or generated with a query $q$.
A.2 Prompts for Explicit Imagine with LLMs
The prompt for explicit imagination of the Imagine Model to imagine a short dummy useful document is:
[…]
Table 7 shows the full prompts for zero-shot results on LLM that we use for open domain QA: NQ, TQA, WQ.
A.3 Experimantal Settings
In this section, we describe the implementation of our experiments in detail, including the baseline methods, backbone models, and hyperparameters. Our model is built based on the T5 Roberts et al. (2020a). Differing from fine-tuning all model parameters $\theta$ of the updated Pre-trained Language Model (LLM), LoRA Hu et al. (2021) freezes all pre-trained Transformer parameters and optimizes only the parameters of each LoRA adapter. We employ LoRA to train a parameter-efficient fine-tuning baseline. Drawing from this, our approach updates only the parameters of the HyperNetwork to generate the weights for each LoRA adapter. This method is adopted based on LongLoRA’s Chen et al. (2023) recommendations and experimental findings, demonstrating improved performance when the normalization and FFN layers components are updated. This is because: 1) dynamically generating LoRA weights enhances generalization and parameter sharing, and 2) LoRA performs comparably to fine-tuning but mitigates the risk of catastrophic forgetting.
For the baseline, most of the hyperparameters are the default parameters of FiD Izacard and Grave (2021). For LoRA Hu et al. (2021), add the LoRA module only to the $\mathcal{QV}$ of the attention layers and also release the normalization and FFN layers.
We consider conducting experiments using three different sizes of T5, namely T5-base, T5-large, T5-3b, and Llama2-7B, Llama2-13B Touvron et al. (2023). Due to memory constraints and online distillation limitations, A100 supports processing 20 documents for T5-3b, while Llama2 does not support distillation. All experiments with T5-3b are conducted on 2 A100 GPUs, T5-large on 2 A6000 GPUs, and T5-Base on 2 RTX 3090 GPUs. However, experiments with Llama2-7b and 13b, except for IMcQA on 2 A100 GPUs, are tested on 8 RTX 3090 GPUs.
A.3.1 Hyperparameters
The detailed hyperparameter setting is as shown in Table 4. For the LoRA modules, we set the $\alpha$ 32 and the lora rank 32.
Models Docu- ments Steps Lr Batch Size
T5 0 40000 1e-4 8
LoRA-Base 0 40000 5e-4 8
IMcQA 0 50000 1e-3 8
LoRA-l 0 40000 1e-4 4
IMcQA-l 0 50000 5e-4 4
FiD-3b 0 40000 1e-4 2
LoRA-3b 0 40000 1e-4 4
IMcQA 0 50000 1e-4 1
LoRA-Base 0$\dagger$ 40000 5e-4 8
IMcQA 0$\dagger$ 50000 1e-3 8
LoRA-l 0$\dagger$ 40000 1e-4 4
IMcQA-l 0$\dagger$ 50000 5e-4 4
LoRA-3b 0$\dagger$ 40000 1e-4 2
IMcQA-3b 0$\dagger$ 50000 1e-4 1
IMcQA 10 50000 5e-4 1
IMcQA-l 10 50000 5e-4 1
FiD-3b 10 40000 1e-4 1
IMcQA-3b 10 50000 1e-4 1
Table 4: Hyperparameter Settings.
Models # Docu- NQ TQA WQ
ments NQ TQA WQ NQ TQA WQ NQ TQA WQ
T5 0 22.16 3.18 4.12 2.65 21.8 3.15 0.88 2.95 28.3
LoRA-Base 0 16.17 4.71 6.89 3.15 21.16 0.00 1.33 3.04 26.38
IMcQA 0 23.89 6.21 10.94 5.31 22.69 6.30 3.23 5.10 30.31
LoRA-Base 1$\dagger$ 37.17 45.20 15.62 19.57 55.37 12.50 14.15 30.89 28.88
IMcQA 1$\dagger$ 40.14 46.61 18.92 24.78 60.75 12.82 17.70 35.24 41.06
FiD 10 46.81 53.93 24.02 28.57 63.32 17.83 18.81 41.88 41.78
IMcQA 10 47.01 55.74 24.13 31.77 64.95 19.52 24.43 48.10 46.36
T5-l 0 28.5* 3.18 4.12 2.65 28.7* 3.15 0.88 2.95 30.6*
LoRA-l 0 17.70 7.49 8.66 3.54 23.87 4.72 0.00 5.65 29.13
IMcQA-l 0 29.32 10.17 14.06 7.02 30.11 7.81 2.65 7.06 32.68
LoRA-l 1$\dagger$ 37.61 48.50 20.71 20.54 62.71 14.81 15.36 33.83 39.37
IMcQA-l 1$\dagger$ 42.32 54.80 22.05 26.11 65.48 18.11 18.58 47.46 45.28
FiD-l 10 46.7* 57.93 25.12 34.29 61.9* 19.64 27.65 53.87 48.1*
IMcQA-l 10 49.92 60.03 25.79 34.35 69.67 20.28 30.19 54.94 51.52
Table 5: OOD results. The primary row in the table header delineates the dataset trained, while the underscored secondary row demonstrates the in-distribution performance. IMcQA attains optimal performance both in-distribution and OOD under diverse document configurations.
A.3.2 Baselines
DPR   Karpukhin et al. (2020) generates by searching for the most relevant documents through dense vector space representation.
FiD   Izacard and Grave (2021) retrieve relevant documents and send them separately to the Encoder, then fuse the information in the Decoder.
GENREAD   Yu et al. (2023) prompt LLMs like InstructGPT Ouyang et al. (2022) to generate a large number of relevant documents and let the reader process them.
LoRA   We use LoRA Hu et al. (2021) to obtain an efficiently fine-tuned baseline and compare it with our method.
A.3.3 Evaluation
For QA datasets, we choose the exact match (EM) score Rajpurkar et al. (2016) as the evaluation metric. An answer is deemed correct if it aligns with any of the responses in the list of acceptable answers after normalization. Normalization involves transforming the text into lowercase, omitting articles, punctuation, and eliminating redundant spaces.
A.4 Downstream Evaluation Datasets
We use the following three Open-Domain QA for the experiments (§ 4.1).
  • NaturalQuestions (Kwiatkowski et al. (2019)) contains questions corresponding to Google search queries. The open-domain version of this dataset is obtained by discarding answers with more than 5 tokens, each accompanied by a Wikipedia article containing the answer.
  • TriviaQA (Joshi et al. (2017)) contains questions gathered from trivia and quiz-league websites. The unfiltered version of TriviaQA is used for open-domain question answering, each question is accompanied by pages from web and Wikipedia searches that may contain the answer.
  • WebQuestions (Berant et al. (2013)) contains questions from web queries matched to corresponding entries in FreeBase Bollacker et al. (2008).
A.5 Full Experimental Results
A.5.1 Supervised Performance
As shown in Table 8, our initial observations indicate that regardless of the method implemented, supplying a certain quantity of related documents can expedite improvement and enhance performance in QA. FiD Izacard and Grave (2021) model outclasses all baseline models in performance. Notably, utilizing FiD-xl with a mere 10 documents yields performance on par with that attained through the use of FiD-l with 100 documents. Larger models not only encapsulate more knowledge but also demonstrate a superior ability to activate and apply this knowledge efficiently.
Additionally, in comparison with LoRA Hu et al. (2021) methods, IMcQA enhances EM scores by an average of +2.2%. In the closed-book setting, the LoRA method manifests a substantial decrease in performance, likely attributable to the inadequacy of learning sufficient knowledge via questions for storage in the LoRA module. On the other hand, IMcQA harnesses both explicit and implicit imaginations to exploit knowledge for improved outcomes. These results indicate that the knowledge stored in the LLMs’ parameters can still be further exploited.
Figure 7: Zero-Shot results (Best_Subspan EM, %) of Llama2-7B on three open-domain QA datasets.
A.5.2 OOD Results
Table 5 shows the full OOD results in QA. It can be observed that our method has the best OOD generalization ability on all three benchmarks. Although LoRA performs well on the in-distribution part, its performance is generally poor on OOD, with some even showing negative performance. This highlights the importance of the domain adaptability of the implicit imagination HyperNetwork in our method, which generates LoRA adapter weights based on input.
A.5.3 Zero-Shot results Best_Subspan EM
LLMs have limited capacity to utilize extensive context effectively and are prone to generating illusions and redundant content. Best_subspan EM assesses whether the answer is included in the output. Previous studies have corroborated that LLMs encapsulate a considerable volume of knowledge and exhibit robust performance in QA.
Here, we report the Best_Subspan_EM values of Llama2-7B and Llama2-13B on three QA datasets. From Figure 7 and Figure 8, it can be observed that Best_Subspan_EM significantly improves, but the EM values are relatively small. This indicates that LLMs may not effectively utilize retrieval documents and are prone to outputting a lot of irrelevant information. Therefore, there is an urgent need to explore efficient techniques that leverage external information and internal knowledge.
However, the model did exhibit a weak adherence to instructions, often failing to output the exact answer. Remarkably, Llama2-13B displayed a decline in EM with an increase in document length on the WQ dataset, whereas the Best_Subspan_EM value augmented. Contrarily, our method excelled in extracting key information by using text imagination during the compression phase.
Figure 8: Zero-Shot results (Best_Subspan EM, %) of Llama2-13B on three open-domain QA datasets.
A.5.4 OOD and Ablation Experiment Results
Models # Docu- Trainable NQ TQA WQ
ments Params NQ TQA WQ NQ TQA WQ NQ TQA WQ
T5 0 220M 22.16 3.18 4.12 2.65 21.8 3.15 0.88 2.95 28.3
LoRA-Base 0 28.3M 5.43 3.15 4.02 0.00 9.60 0.00 0.22 1.77 20.47
         w FFN 0 141.5M 16.17 4.71 6.89 3.15 21.16 0.00 1.33 3.04 26.38
         w FFN & LCD 0 141.5M 21.37 2.82 6.89 1.99 17.94 3.74 0.00 2.82 32.50
IMcQA 0 26.1M 5.31 3.82 5.71 0.22 10.34 2.12 0.55 2.30 16.58
         w FFN 0 139.3M 21.05 4.52 6.50 3.51 19.08 3.15 2.11 3.84 28.17
         w FFN & LCD 0 141.5M 23.89 6.21 10.94 5.31 22.69 6.30 3.23 5.10 30.31
T5-l 0 770M 28.5* 3.18 4.12 2.65 28.7* 3.15 0.88 2.95 30.6*
LoRA-l 0 42.5M 4.42 6.50 7.87 3.98 10.03 3.94 1.99 6.71 18.11
         w FFN 0 445.1M 17.70 7.49 8.66 3.54 23.87 4.72 0.00 5.65 29.13
         w FFN & LCD 0 445.1M 28.32 4.52 10.94 5.31 25.71 6.12 1.75 4.52 29.92
IMcQA-l 0 34.8M 7.08 8.90 9.45 4.42 13.14 8.66 2.43 10.17 17.72
         w FFN 0 437.5M 23.01 8.33 11.02 3.51 20.08 3.15 3.51 5.65 31.50
         w FFN & LCD 0 437.5M 29.32 10.17 14.06 7.02 30.11 7.81 2.65 7.06 32.68
Table 6: OOD and ablation experiment results. * denotes the results are from the existing papers and LCD denotes Long Context Distillation.
Here, we supplement the experimental results of LoRA and IMcQA under supervised fine-tuning in closed-book settings and the ablation results of feedforward neural network (FFN) and Long Context Distillation (LCD). It can be observed that our method like LoRA, belongs to parameter-efficient fine-tuning, and because we share the HyperNetwork to generate LoRA adapter weights, we fine-tune fewer parameters.
From Table 6, it can be seen that releasing FFN can bring more performance improvement, possibly because adding LoRA in Attention cannot fully utilize enough knowledge Yao et al. (2022). With the support of LCD, performance is further improved, with an average increase in EM of +5%. This also proves the effectiveness of our proposed LCD. In comparison with IMcQA and LoRA, it becomes more evident that LoRA tends to transfer knowledge to the LoRA module, resulting in low generalization. Our method enhances knowledge activation through dynamic generation, showing significant effects not only in-distribution but also in OOD.
Methods Prompt
CBQA Please write a high-quality answer for the given question using your knowledge. Only give me the answer and do not output any other words. Question: {question} Answer:
Retrieval Please write a high-quality answer for the given question using only the provided search results (some of which might be irrelevant). Only give me the answer and do not output any other words. Context: {context} Answer the question based on the given passages. Question: {question} Answer:
Imagine Please write a high-quality answer for the given question using your knowledge and the provided imagined compressed results (some of which might be irrelevant). Only give me the answer and do not output any other words. Imagined Context: {context} Answer the question based on your knowledge and the given imagined context. Question: {question} Answer:
Table 7: Prompts for different methods on Zero-Shot setting. CBQA denotes closed-book QA that just prompts the model with the question.
Models Reader Params # Documents NQ TriviaQA WebQ
# closed-book setting
T5 Roberts et al. (2020a) 220M 0 25.9* 23.8* 27.9*
T5-l Roberts et al. (2020a) 770M 0 28.5* 28.7* 30.6*
T5-xl 3b 0 28.30 33.92 34.43
LoRA-Base Hu et al. (2021) 220M 0 16.17 21.16 26.38
LoRA-l Hu et al. (2021) 770M 0 17.70 23.87 29.13
LoRA-xl 3b 0 23.15 32.16 35.24
IMcQA (Ours) 220M 0 23.89 22.69 30.31
IMcQA-l (Ours) 770M 0 29.32 30.11 32.68
IMcQA-xl (Ours) 3b 0 29.59 35.71 37.40
# Retrieval Augmented Generation
DPR Karpukhin et al. (2020) 110M 100 41.5* 56.8* 41.1*
RAG Lewis et al. (2020) 400M 10 44.5* 56.1* 45.2*
FiD Izacard and Grave (2021) 220M 100 48.2* 65.0* 46.71
FiD-l 770M 100 51.4* 67.6* 50.52
FiD-xl 3b 20 55.18 72.92 52.85
FiD-l 770M 10 46.7* 61.9* 48.1*
FiD-xl 3b 10 50.1* 66.3* 50.8*
# Generation Augmented Generation
GENREAD-l Yu et al. (2023) 770M 10$\dagger$ 40.3* 67.8* 51.5*
GENREAD-xl 3b 10$\dagger$ 42.6* 69.6* 52.6*
# Our proposed method
LoRA-Base 220M 1$\dagger$ 34.51 54.05 32.28
LoRA-l 770M 1$\dagger$ 40.05 62.81 43.70
LoRA-xl 3b 1$\dagger$ 44.15 66.92 48.23
IMcQA 220M 1$\dagger$ 40.14 60.75 41.73
IMcQA-l 770M 1$\dagger$ 42.32 65.48 45.28
IMcQA-xl 3b 1$\dagger$ 46.51 68.38 50.45
IMcQA 220M 10 47.01 64.95 46.36
IMcQA-l 770M 10 49.92 69.67 51.52
IMcQA-xl 3b 5$\ddagger$ 50.87 70.34 52.78
Table 8: Full QA performances (%) of different methods on three datasets. The first part (closed-book setting) indicates that explicit documentation was not utilized; The latter three parts utilize explicit augmented documents. The best results are in bold. * means that those results are from existing papers, $\dagger$ denotes that the number of documents is generated ($\ddagger$ indicates that the number of documents is reduced due to insufficient memory for distillation).