### v3

# Imagination Augmented Generation: Learning to Imagine Richer Context for Question Answering over Large Language Models

Abstract
Retrieval-Augmented-Generation and Gener-ation-Augmented-Generation have been proposed to enhance the knowledge required for question answering over Large Language Models (LLMs). However, the former relies on external resources, and both require incorporating explicit documents into the context, which increases execution costs and susceptibility to noise data. Recent works indicate that LLMs have modeled rich knowledge, albeit not effectively triggered or awakened. Inspired by this, we propose a novel knowledge-augmented framework, Imagination-Augmented-Generation (IAG), which simulates the human capacity to compensate for knowledge deficits while answering questions solely through imagination, thereby awakening relevant knowledge in LLMs without relying on external resources. Guided by IAG, we propose an imagine richer context method for question answering (IMcQA). IMcQA consists of two modules: explicit imagination, which generates a short dummy document by learning from long context compression, and implicit imagination, which creates flexible adapters by distilling from a teacher model with a long context. Experimental results on three datasets demonstrate that IMcQA exhibits significant advantages in both open-domain and closed-book settings, as well as in out-of-distribution generalization111 Our code will be available at https://anonymous.4open.science/r/IMcQA.
1 Introduction
Knowledge-intensive tasks like question answering (QA) necessitate utilizing extensive world and domain knowledge Berant et al. (2013); Joshi et al. (2017); Kwiatkowski et al. (2019). Nowadays, Large Language Models (LLMs) have displayed notable competencies in almost every task and industry Liu et al. (2023b). However, LLMs lack the sufficient capability to independently handle knowledge-intensive tasks Frisoni et al. (2024) and usually generate hallucinations Zhao et al. (2023).
Figure 1: Compared with RAG and GAG , the proposed IAG eschews external resources, generates a shorter context (explicitly imagination) and creates flexible adapters (implicitly imagination) for each question.
In recent years, to mitigate hallucinations in LLMs and improve performance in knowledge-intensive tasks such as QA, researchers have proposed numerous knowledge-augmented methods for LLMs. These methods primarily fall into two categories: Retrieval-Augmented-Generation (RAG) Guu et al. (2020) and Generation-Augmented-Generation (GAG) Kim et al. (2024). RAG (Top part of Figure 1) retrieves related documents from external resources (e.g., Wikipedia) and then inputs both the retrieved documents and the question into LLMs Izacard and Grave (2021). GAG (Middle part of Figure 1) leverages external LLMs such as ChatGPT Ouyang et al. (2022) to generate relevant documents Yu et al. (2023).
However, these methods have the following disadvantages [^2] A more intuitive comparison can be seen in A.1.: 1) Dependence on external resources, RAG relies on external domain knowledge resources Ke et al. (2024), while GAG depends on a more powerful external LLM as a knowledge generator. This dependency hinders their broader application. 2) Increased execution costs, The required computing resources and inference time continue to increase as the number of documents increases Xu et al. (2023a). For example, the typical RAG method FiD Izacard and Grave (2021) needs to handle over 12K tokens for retrieving 100 documents, leading to more than a $100$-fold increase in prompt length and over $100^{2}$-fold increase in inference time Liu et al. (2023a). The GAG method Yu et al. (2023) requires additional financial costs (e.g., API calls). 3) Susceptibility to noise data, both methods are prone to the inclusion of irrelevant or noisy data Li et al. (2023); Shaier et al. (2024). For example, Gao et al. (2024) indicates that noise in the documents negatively affects the performance. Therefore, there is an urgent need to explore new knowledge-augmented methods.
In fact, LLMs inherently contain rich knowledge and possess significant potential for resolving knowledge-intensive tasks Bhagavatula et al. (2020). Enhancing the performance of specific tasks can be achieved by better activating and awakening relevant knowledge or expanding memory capacity without relying on external resources. For example, simply repeating the question twice Xu et al. (2023b), just reviewing and consolidating knowledge by appending a straightforward prompt “As far as I know” Yao et al. (2023), and using visual-language models to imagine images Tang et al. (2023), all these approaches enhance the performance of LLMs on downstream tasks.
Therefore, inspired by the above phenomena and aiming to address the challenges faced by RAG and GAG, we propose a novel knowledge-augmented framework, Imagination-Augmented-Generation (IAG). IAG simulates the human capacity to compensate for knowledge deficits solely through imagination in QA. For resolving knowledge-intensive tasks, IAG utilizes solely LLMs to generate a shorter context and create a more flexible adapter (Bottom part of Figure 1). Through a knowledge distillation framework, IAG achieves knowledge enhancement through imagination. The fundamental idea is to permit student models, which are deficient in rich contextual information (e.g., without abundant retrieved documents), to imitate teacher models that are endowed with such information.
Within the framework of IAG, we introduce an imagine richer context method for question answering (IMcQA). To sufficiently utilize the inherent knowledge of LLMs, we design two main modules to activate the various potential knowledge modeling in LLMs and obtain a richer context. Specifically, the explicit imagination module first uses symbol distillation to obtain the compressed context and then guides the context generator to generate a short dummy document. Subsequently, within the knowledge distillation framework, the implicit imagination module utilizes the hypernetwork to generate adapters to awaken the question-aware ability of LLMs. Unlike the LoRA Hu et al. (2021) stores task knowledge and ability in modules, the hypernetwork learns to generate a question-aware knowledge module.
We evaluate the proposed IMcQA on various LLMs, including T5 Roberts et al. (2020a) and Llama2 Touvron et al. (2023). The experimental results across three QA datasets indicate that the proposed method yields performance gains while reducing computational expenses and time. Notably, it even outperforms baseline methods that retrieve and generate knowledge under the same document settings. In conclusion, the contributions of this paper are summarized as follows:
  • We propose a new knowledge augmentation framework IAG to fully leverage the LLMs’ intrinsic knowledge more efficiently without relying on external resources.
  • We propose a novel QA method IMcQA that employs two main modules (explicit imagination and implicit imagination) to utilize better the knowledge stored in the LLMs and awaken a richer context in QA.
  • Experimental results indicate that the proposed method successfully activates the relevant internal knowledge of LLMs. IMcQA exhibits significant advantages in both open-domain and closed-book settings, as well as in out-of-distribution generalizations.
2 Related Work
This paper mainly utilizes context compression, hypernetworks and knowledge distillation to achieve knowledge enhancement. The following will elucidate pertinent research across four facets.
Knowledge Enhancement has usually been adopted to alleviate the issue of insufficient knowledge in LLMs. There are two main methods: RAG Sun et al. (2019); Wang et al. (2024) and GAG Abdallah and Jatowt (2023). The typical RAG method FiD Izacard and Grave (2021) retrieves documents from Wikipedia to answer questions. LLMs serving as a knowledge base have been the focus of numerous studies that advocate the extraction of knowledge from such models (e.g., GPT-3). For instance, Yu et al. (2023) generates 10 documents for each question. However, RAG requires external resources, and both RAG and GAG need verbose long contexts. Recently, methods have been developed to enhance LLMs’ abilities by simulating human imagination of visual information using existing visual-language models Tang et al. (2023); Akter et al. (2024). We prefer self-imagination to augment knowledge and aim to leverage the parameterized knowledge within the models Xu et al. (2023b); Kazemnejad et al. (2023). Our proposed method not only eliminates the need for external resources but also improves the efficiency of activating internal knowledge within LLMs.
Context Compression has often been used to improve the efficiency of LLMs in processing long contexts. Recent studies Mu et al. (2023) propose that long contexts be condensed into summary vectors (soft prompts) to ensure their effective utilization by LLMs. Simultaneously, some studies Jiang et al. (2023); Pan et al. (2024) suggest utilizing information redundancy and entropy in lengthy texts to compress contexts Li et al. (2023). Unlike these approaches, this paper aims to enhance the long-context modeling ability of LLMs. By developing a context generator that creates compressed contexts, the QA model operating on short contexts can achieve a rich contextual understanding similar to models designed for longer contexts.
Knowledge Distillation is a technique where a smaller model learns to mimic the predictions of a larger model, aiming to retain performance while reducing computational resources Hinton et al. (2015). Recent studies West et al. (2022) present symbolic knowledge distillation, a process that facilitates knowledge transfer from a teacher model via extracting training data to subsequently train a student model Wang et al. (2023b); Ranaldi and Freitas (2024). In this paper, the process of obtaining compressed context during the explicit imagination resembles a form of symbolic distillation. Regarding training, our emphasis lies in distilling the long-context modeling abilities of LLMs.
Hypernetworks is designed to reduce the number of parameters Ha et al. (2016), i.e., a small neural network generates parameters for another big neural network. Recent studies Phang et al. (2022); Ivison et al. (2023) have explored the enhancement of model performance in zero/few-shot settings through meta-learning involving hypernetworks. We utilize hypernetworks to acquire implicit imagine capabilities by dynamically generating LoRA for efficiency and generalization.
3 Method
In this section, we introduce the detailed method of IMcQA to activate LLMs’ intrinsic knowledge and obtain a richer context for QA. The fundamental premise underlying this method is that QA with a richer context yields greater performance (e.g., RAG with retrieved documents). Consequently, diverse methods are employed for questions lacking in richer contexts to activate knowledge within LLMs to replicate comparable effects to those achieved with richer contexts.
Specifically, IMcQA comprises two main modules. Explicit imagination with long context compression learns to imagine a short dummy document (§ 3.2). And implicit imagination with the hypernetwork models’ hidden knowledge that learns a shared knowledge feature projection across questions (§ 3.3). The hypernetwork is trained to generate lightweight LoRA modules, aiming to align the question and the internal knowledge. Besides, there is long context distillation in training, which learns the teacher’s rich representations to compensate for missing knowledge in imagination (§ 3.4).
3.1 Formulation
The formulation of our task follows RAG for QA Guu et al. (2020). Let $\mathcal{V}^{*}$ denote the infinite set, encompassing all potential strings over the tokens in vocabulary $\mathcal{V}$, and this includes the empty string. An instance within a QA dataset is defined as a triplet $(q,a,c)$ comprising question $q$, answer $a$, and context $c$, where $q,a,c\in\mathcal{V}^{*}$. Conventionally, the context $c$ is drawn from the knowledge corpus $\mathcal{Z}$, like Wikipedia, whereby $\mathcal{Z}\subset\mathcal{V}^{*}$. Additional background details are available in B.1.
Figure 2: Overview of IMcQA method. In the inference phase, for each question, the explicit imagination generates a short dummy document and the implicit imagination with hypernetwork creates a specific LoRA module. During training, there are two stages: the first stage is the pre-training of the context generator (§ 3.2), aiming at its ability to imagine a short dummy document based on the question, and the second stage is the hypernetwork fine-tuning (§ 3.3) using long context distillation (§ 3.4) to obtain a question-specific LoRA module.
3.2 Explicit Imagination with Context Generator
To obtain a context $c$ for each question, we utilize LLMs to generate a short dummy document, which can mitigate knowledge corpus error Lee et al. (2023) by considering potentially useful contexts. We employ the "teacher-student" framework. During the training phase, the teacher model, equipped with rich contextual information, generates short contexts through context compression. The student model, in turn, learns to automatically generate these short contexts by leveraging its own LLM capabilities.
As shown in the left part of Figure 2, to help LLMs fully utilize the knowledge and generate compressed text, we first pretrain the context generator on our collected question-compressed document pairs. By leveraging symbolic distillation, we employ the long context compression method LongLLMLingua Jiang et al. (2023) to condense a large corpus of retrieved documents. These compressed texts $c^{\prime}$ then serve as fine-tuning data alongside specific prompts $p_{q}$ (B.2) and question-answer pairs for the context generator [^3] We discuss the role of context generator in the A.2. $G_{\theta}$ ($\theta$ represents the model’s parameter), which guides the model to think about its knowledge and generate a short dummy document:
$d=G_{\theta}(p_{q}(q;c^{\prime}))$ (1)
where $d$ is the dummy document generated from the context generator. This process enables LLMs to conceive compressed knowledge that robustly parallels the question’s knowledge requirements.
3.3 Implicit Imagination with Hypernetwork
Generally speaking, richer context can help LLM better answer questions. That is, the representation of questions and the internal state of LLM when utilizing rich context are the better states. Therefore, in the absence of context, we should focus on building models to awaken LLM to achieve this better state and as a better QA model.
We utilize the hypernetwork [^4] We conduct a detailed analysis of the reasons behind the hypernetwork in the A.3. to learn to generate a better QA model, which does not directly optimize the LoRA Hu et al. (2021) module but generates specific LoRA adapter weights using the inputs for QA (bottom part of Figure 2). This is akin to repeating the question in the prompt Xu et al. (2023b) and incorporating certain topical cues to stimulate the model’s recall of relevant questions Wang et al. (2023c). However, the distinction lies in the fact that they serve as wake-up features, whereas we are generating model parameters.
Figure 3: The Architecture of hypernetwork. Hypernetwork generates LoRA adapter weights for each question. During training, only Hypernetwork, FFN, and Norm weights are updated.
The hypernetwork architecture for generating LoRA weights is exhaustively outlined in Figure 3. Specifically, $D_{k}^{q}$ and $U_{k}^{q}$ represent the low-rank projections of layer $k$ correlated with the $\mathcal{Q}$, while $D_{k}^{v}$ and $U_{k}^{v}$ correspond to those associated with the $\mathcal{V}$. Hypernetwork represented as $g_{D}$ and $g_{U}$, takes $\textit{concat}(f,i_{k}^{\{q,v\}})$ as input, where $f$ is the feature vectors that use the model’s encoder to obtain and using whitening algorithm Su (2021) for dimensionality reduction, $i_{k}^{\{q,v\}}\in\{0,\ldots,2\times\#\text{blocks}\}$ signifies the positional embedding differentiating between layers and between $\mathcal{QV}$. Each hypernetwork is defined by weights $W_{d}$ and $W_{u}$ which represent the down and up projections respectively. Finally, the hypernetwork equations for $D^{\{q,v\}}$ and $U^{\{q,v\}}$ can be expressed as:
$f=\text{whitening}(\text{Encoder}(q;d))$ (2)
$x=\text{concat}(f,i_{k}^{\{q,v\}}~{}|~{}i_{k}^{q}=2k,i_{k}^{v}=2k+1)$ (3)
$D^{\{q,v\}},U^{\{q,v\}}=g_{D}(x),g_{U}(x)$ (4)
where the Encoder represents the encoder of the model, whitening is a dimensionality reduction algorithm, and concat means to splice the content. $g_{D}$ and $g_{U}$ denote the descending and ascending dimensions of hypernetwork. More formally,
$g(x)=\text{MM}(\text{ReLU}(\text{MM},W_{d}),W_{u})$ (5)
where MM stands for matrix multiplication, ReLU is a activation function.
3.4 Training with Long Context Distillation
Within the framework of knowledge distillation, components such as hidden representations Jiao et al. (2020), attention dependencies Wang et al. (2020), and relations among representations Park et al. (2021) are regarded as valuable knowledge for transfer. In this paper, we consider long context distillation (LCD) as the contextualized knowledge that mainly guides the student.
Specifically, the teacher model FiD Izacard and Grave (2021), which utilizes longer contextual inputs and theoretically contains more information (richer context). It will activate more specific internal knowledge and serve as a supervisory model. The teacher model assists the student model T5 Roberts et al. (2020a), which has the same size as the teacher and leverages short contextual inputs. This aids in activating richer feature representations and knowledge. The optimization objective for the student model at each mini-batch $z_{r}=(x_{r},y_{r})$ is:
$\begin{split}\mathcal{L}_{\text{s}}(\theta_{s},\theta_{t},z_{r})=\alpha\mathcal{L}_{\text{ce}}(y_{r},S(x_{r};\theta_{s}))\\ +(1-\alpha)\mathcal{L}_{\text{ce}}(T(x_{r};\theta_{t}),S(x_{r};\theta_{s}))\end{split}$ (6)
where we have a teacher model denoted as $T(\cdot;\theta_{t})$ and a student model denoted as $S(\cdot;\theta_{s})$. The corresponding model parameters are $\theta_{t}$ and $\theta_{s}$.
As shown in the right of Figure 2, we perform an additional representation alignment for better knowledge transfer. In our distillation, both teacher model and student model have the $L$ layers, we feed the text into them and can obtain the corresponding output hidden states $\{H_{l}^{t}\}_{l=0}^{L}$, $\{H_{l}^{s}\}_{l=0}^{L}$, and attention matrices $\{A_{l}^{t}\}_{l=1}^{L}$, $\{A_{l}^{s}\}_{l=1}^{L}$. We suppose the student’s $l$-th layer is aligned with the teacher’s $l$-th layer, then the outputs of the student (i.e., $H_{l}^{s}$ and $A_{l}^{s}$) should be close to the teacher’s (i.e., $H_{l}^{t}$ and $A_{l}^{t}$). For aligning hidden states, following Park et al. (2021), we use cosine distance COS to calculate the proximity between the hidden states of the teacher and the student:
$\mathcal{L}_{\text{hid}}=-~{}\text{COS}(H_{l}^{s},H_{l}^{t})$ (7)
While for aligning attention dependencies, we follow Jiao et al. (2020) to optimize the mean square error (MSE) between the attention matrices of the teacher and the student:
$\mathcal{L}_{\text{attn}}=-~{}\text{MSE}(A_{l}^{s},A_{l}^{t})$ (8)
The overall objective for knowledge transfer is:
$\mathcal{L}_{\text{align}}(H_{l}^{s},H_{l}^{t},A_{l}^{s},A_{l}^{t})=\mathcal{L}_{\text{attn}}+\mathcal{L}_{\text{hid}}$ (9)
The overall objective for training IMcQA is the weighted sum of the two objectives:
$\mathcal{L}=\mathcal{L}_{\text{s}}+\lambda\mathcal{L}_{\text{align}}$ (10)
Models # Docs NQ TriviaQA WebQ
Large (800M) XL (3B) Large (800M) XL (3B) Large (800M) XL (3B)
# Closed-book Setting
T5 Roberts et al. (2020a) 0 28.5 28.30 28.7 33.92 30.6 34.43
LoRA Hu et al. (2021) 0 17.70 23.15 23.87 32.16 29.13 35.24
IMcQA (Ours) 0 29.32 29.59 30.11 35.71 32.68 37.40
# Retrieval Augmented Setting (compared with RAG)
DPR Karpukhin et al. (2020) (110M) 100 41.5 - 56.8 - 41.1 -
RAG Lewis et al. (2020) 10 44.5 - 56.1 - 45.2 -
FiD Izacard and Grave (2021) 10 46.7 50.1 61.9 66.3 48.1 50.8
FiD Izacard and Grave (2021) 100 51.4 55.2 67.6 72.9 50.5 52.9
EAR Chuang et al. (2023) 10 39.6 42.3 60.0 64.6 - -
RFiD Wang et al. (2023a) 10 48.3 50.5 63.4 67.8 - -
FILCO Wang et al. (2023d) 1 - 44.7 - 59.0 - -
IMcQA (Ours) 10 49.9 50.9 69.7 70.3 51.5 52.8
IMcQA (Ours) 30 53.1 - 70.5 - 52.0 -
# Generation Augmented Setting (compared with GAG)
GENREAD (sampling) Yu et al. (2023) 10 40.3 42.6 67.8 69.6 51.5 52.6
GENREAD (clustering) Yu et al. (2023) 10 43.5 45.6 70.2 71.6 53.5 54.4
IMcQA (Ours) 10 48.8 49.2 70.9 72.2 54.5 55.6
# Imagination Augmented Setting
LoRA Hu et al. (2021) 1 40.1 44.2 62.8 66.9 43.7 48.2
IMcQA (Ours) 1 42.3 46.5 65.5 68.4 45.3 50.5
Table 1: QA performances of different methods with different settings. The first part (closed-book setting) indicates that only utilize questions; The latter three parts utilize explicit documents. The best results are in bold, while the second-best are underlined. * means that those results are from existing papers, denotes that the documents were generated ( indicates that the number of documents is reduced due to insufficient memory for distillation).
4 Experiment
In this section, we conduct experiments to demonstrate the effectiveness and efficiency of IAG and IMcQA on QA. The experiment mainly answers four research questions (RQs):
RQ1: Can IAG and IMcQA achieve knowledge augmentation for QA over LLMs? (§ 4.4)
RQ2: Does IMcQA have a good out-of-distribution generalization ability? (§ 4.5)
RQ3: Does IAG have advantages in effectiveness and efficiency compared to RAG and GAG? (§ 4.6)
RQ4: What is the role of explicit and implicit imagination modules in IMcQA? (§ 4.7)
4.1 Datasets
We evaluate the proposed approach on three public question answering datasets: NaturalQuestions (NQ) Kwiatkowski et al. (2019), WebQuestions (WQ) Berant et al. (2013) and TriviaQA (TQA) Joshi et al. (2017). To evaluate the model performance, we use the exact match (EM) score for evaluating predicted answers Rajpurkar et al. (2016). We provide dataset details in the B.4.
4.2 Baselines
[…]
To demonstrate the plug-and-play capability of IMcQA on the zero-shot settings of LLMs ($\geq$ 3B), we use Llama2-7B and -13B Touvron et al. (2023) as the basic model. We evaluate with 6 diverse settings: without retrieval, with retrieval, with LoRA, RECITE Sun et al. (2023), HICL Wang et al. (2024) and using the proposed IMcQA.
4.3 Implementations
In the pretraining stage, the context generator initialized with T5-large utilizes the generated question-compressed pairs. During the second stage, the teacher model employs a FiD reader with different sizes (FiD-l and FiD-xl) that are fine-tuned on the training split of target datasets. The student model freezes the backbone and updates solely the hypernetwork [^5] B.3 contains more implementation and baseline details..
4.4 Main Results
4.4.1 Supervised Setting
Table 1 shows the performance results, full results including T5-Base are in the C.1. When juxtaposed with closed-book models, RAG, and GAG methods, our proposed IAG framework IMcQA method exhibits state-of-the-art (SOTA) performance with the equivalent magnitude of document numbers.
In the closed-book setting (in the upper part of the table), our method outperforms the baseline by an average of +2% EM score, indicating its excellence in utilizing internal knowledge with imagination. It’s especially noteworthy that as the model size expands, the performance advantages of the imagination become ever more evident.
The following three parts show the experimental results under the open domain setting [^6] Due to memory constraints from online distillation, we only compare the IMcQA-l under the RAG setting using 30 documents with FiD using 100 documents.. Although our method only deals with one short dummy document, it can still achieve results similar to or better than the RAG and GAG methods, which handle 10 documents. The findings reveal that IMcQA exploits imagined condensed text to strike a balance between efficiency and overhead. Beyond this, when IMcQA implements 10 retrieved documents under RAG, it overtakes RFiD’s performance, posting 1.6% higher in NQ, 4.4% in TQA, and 2.7% in WQ. IMcQA employing 10 generated documents under GAG outstrips GENREAD (clustering)’s performance, documenting a 4.5% increment in NQ, 0.7% in TQA, and 1.1% in WQ.
Figure 4: Zero-Shot results (EM, %) of Llama2-7B on three open-domain QA datasets. The number in parentheses indicates the number of documents used.
4.4.2 Zero-shot Setting
Figure 4 illustrates the zero-shot results for LLMs implementing IMcQA with a frozen Llama2-7B and -13B. This research seeks to explore the possibility of enhancing LLMs via IAG. Due to the high computational demands of training, we only fine-tuned the hypernetwork on a mixed dataset without LCD in this experiment and evaluated performance in a zero-shot setting [^7] Detailed prompts information can be found in the B.2.. More zero-shot setting results can be seen in C.3.
We discerned that Llama2’s performance can be enhanced by imagining knowledge autonomously. While leveraging explicit imagined context could amplify the average EM +1%, this is not as significant as the improvement achieved by retrieving 10 documents, indicating the limitations of relying solely on prompt cues for triggering corresponding knowledge. IMcQA can enhance knowledge via two main imagination processes, escalating EM by +15.33% for NQ, +11.97% for TQA, and +16.38% for WQ. Compared to two other advanced RAG methods, IMcQA using a single document, performs only 1 EM lower than the HICL method Wang et al. (2024) on the TQA but achieves +10% EM on the NQ and +5% EM on the WQ. With IAG, Llama2-7B demonstrated an average improvement of +14% across the three datasets. This trend is also observed in Llama2-13B’s results (Figure 6). This implies that even in zero-shot settings, our method can still offer substantial benefits to LLMs.
Models # Docs Base (220M) Large (800M)
NQ TQA WQ NQ TQA WQ
T5 0 22.16 3.18 4.12 28.5* 3.18 4.12
IMcQA 0 23.89 6.21 10.94 29.32 10.17 14.06
FiD 10 46.81 53.93 24.02 46.7* 57.93 25.12
IMcQA 10 47.01 55.74 24.13 49.92 60.03 25.79
LoRA 1 5.43 3.15 4.02 4.42 6.50 7.87
IMcQA 1 40.14 46.61 18.92 42.32 54.80 22.05
Table 2: IID and OOD results. The performance on three open-domain datasets for the model trained on NQ is reported, with the underlined values indicating the IID performance. Full OOD results and details of three datasets are shown in the C.2.
4.5 Out-Of-Distribution (OOD) Performance
To further demonstrate the generalizations of the IMcQA method and the importance of hypernetwork, we also evaluate its performance in OOD generalizations. Table 2 shows the IID and OOD performance of FiD, and IMcQA methods with different document settings when training on NQ (From NQ generalization to the other two datasets).
It is patently clear that an increment in document provision leads to better OOD performance, likely due to the presence of answer-oriented content within these documents. Remarkably, IMcQA can come within a relatively narrow 5% gap of FiD, even when utilizing a single imagined document as opposed to 10 retrieved documents.
Simultaneously, IMcQA generally showcases superior performance in OOD when provided with 10 retrieved documents. This superiority can be traced back to the pivotal role played by hypernetwork in generating LoRA adapters’ weights based on questions. This equips models with the capability to invoke and access internal knowledge based on context-specific discourse rather than confining to resolving distinct questions.
Models Training Params # Documents # Avg Tokens Inference Time GPU Memory Training Time
T5 220M 0 19.8 79.8s 2828M 0.9h
IMcQA 139.3M 0 19.8 82.3s 2710M 1.2h
IMcQA 139.3M 1 522.1 214.6s 2882M 1.7h
FiD 220M 10 1748.3 683.3s 4358M 2.3h
GENREAD 220M 10 1912.5 704.8s 4412M -
FiD 220M 100 16625.7 1293.2s 19048M 5.8h
Table 3: Training and inference cost on the NQ. The backbone model is T5-Base.
4.6 Training Cost and Inference Speed-up
We proceeded to measure the inference speed documented in GPU time and training time for 5000 steps on the NQ dataset using T5-Base. The experiments were conducted on a single RTX 3090 GPU, maintaining a standard batch size of 8 during training and 1 during inference.
As evident from Table 3, the proposed method’s advantage lies in its diminished requirement for parameter updates, which can be attributed to the shared hypernetwork’s utilization that generates LoRA adapters, thereby negating the necessity of individual LoRA adapters’ setup. Despite the lack of a training advantage due to distillation constraints, IMcQA achieves efficient reasoning through an extremely lightweight design, saving more than half the training time compared to methods using a large number of documents (×$\times 0.3$). Compared to the other two methods, the processed tokens are significantly decreased, while either outperforming them or showing negligible differences in performance. This represents an optimal trade-off between efficiency and computational demand. Moreover, unlike GAG, our approach incurs no financial costs associated with API calls, and the reduced model size facilitates faster generation.
4.7 Ablation Experiment
This study introduces two key imagination processes to stimulate LLMs’ internal knowledge: Explicit Imagination (EI) and Implicit Imagination (II). We particularly examined the influence of different imagination types on performance.
Figure 5 demonstrates that both EI and II are important for IMcQA. Omitting either one results in a considerable reduction in performance, with a drop exceeding 10% observed when EI is neglected. This is harmonious with the initial observation that performance improvement becomes more noticeable when relevant documents are available, thus underscoring EI’s superiority.
The outcomes of Long Context Distillation (LCD) and the application of EI in the hypernetwork also make marginal contributions to the overall results. This validates the previous assertion that a more extensive context tends to optimize performance, although with limited gains. The experiments and analysis above demonstrate the importance of each component and the effectiveness of our IMcQA method.
Figure 5: Ablation experiment results (%) of T5-Base on three open-domain QA datasets.
5 Conclusion and Future Work
This study proposes a novel knowledge-augmented strategy for LLMs, namely Imagine Augmented Generation (IAG), and a specific method IMcQA for open domain question answering. The proposed method effectively activates and utilizes intrinsic knowledge within LLMs through two imaginations: explicit imagination, and implicit imagination. Experimental results demonstrate a significant improvement in QA performance while remaining relatively lightweight. Although the main focus of this method is on one specific task, we believe these findings can offer a novel perspective on how to better harness the potential of LLMs. In the future, we plan to apply IAG to more NLP tasks and explore multimodal knowledge-augmented generation.
Limitations
While this study has demonstrated significant achievements in QA tasks, there are notable limitations:
Tasks.   The proposed methods in the study are specialized specifically for QA. It remains unknown how effective they would be in other types of knowledge-intensive tasks, such as fact-checking or dialogue systems. Further validation is needed to assess the generalizations and applicability of this approach.
Multimodal.   We have only considered imagined text and hidden representations. In future work, it is imperative to explore multimodal information including the impact of imagining images on performance.
Method.   Our method relies on the knowledge learned by LLMs in the pre-training phase, which may limit the model’s ability to quickly adapt to new information. The dependency on internal knowledge activation in IAG may lead to a less transparent decision-making process in the model, making it challenging to explain the logic behind the generated answers. In the future, there is a need to continue exploring adaptive knowledge enhancement methods to optimize results further.
Ethical Considerations
In this paper, we proposed a novel knowledge enhancement method aimed at leveraging the knowledge of LLMs. However, LLMs may generate inappropriate or discriminatory knowledge. Our approach does not introduce ethical concerns. The datasets we used are public, and there are no privacy issues.
Appendix A Method
A.1 Comparison of Three Paradigms
Compared to RAG and GAG, our method has certain limitations, such as requiring a more complex training process and the necessity of training a model. Similar to the GAG method, which uses a master’s degree in law as a knowledge base, our method also struggles to generate content when encountering new and unknown world knowledge, which presents a challenge that needs to be addressed. Additionally, the knowledge base might be affected by knowledge gaps in low-resource settings where there is a lack of a comprehensive knowledge base.
Next, we compare IAG, RAG, and GAG across four criteria for a more intuitive understanding. From the table 4, it can be observed that the document relevance obtained by IAG and GAG is higher, while RAG heavily relies on the retriever and external knowledge base. In terms of document length usage, IAG only needs to use a virtual document, greatly reducing the number of tokens. Therefore, IAG is superior to the other two methods in terms of reasoning time.
A.2 Context Generator
There are two main goals in the pre-training of the model in the first stage of IAG (context generator): first, to improve its ability as a document generator by learning to generate rich and concise documents; second, to introduce some external knowledge that the model initially does not possess. It’s worth noting that the second goal is crucial, as the model may encounter knowledge it has not yet learned. Thus, IAG does not rely on external large models or retrievers for external reasoning and can complete reasoning independently.
A.3 Hypernetwork
Hypernetworks have gained significant attention in recent years due to their potential to enhance various aspects of neural network performance. In this section, we analyze the reasons for employing hypernetworks in detail:
Hypernetworks, which are neural networks designed to generate the weights of other networks, allow for dynamic adjustment of model parameters. This adaptability enables the model to better suit different tasks and datasets, thereby improving overall performance. By utilizing hypernetworks, the number of models that need to be trained individually can be significantly reduced. Traditional methods require separate models for each task, whereas hypernetworks can generate weights for multiple tasks. This capability enhances training efficiency. In our task, we use hypernetworks to generate adapters for the question and input, which are then inserted into the model. This helps the model incorporate the knowledge targeted by the question, corresponding to implicit imagination. Compared to traditional efficient fine-tuning, this process is more aligned with the goal of imagination.
Hypernetworks can capture the commonalities and differences between various tasks by learning to generate weights. This ability to generalize across tasks improves the model’s performance on unseen data, making it more robust in diverse scenarios. In multi-task learning or meta-learning scenarios, hypernetworks can considerably reduce the need for storing multiple independent models. A hypernetwork only needs to store a single generating network and some shared parameters, thus significantly decreasing the storage space required. Hypernetworks can quickly generate new weights to adapt to new tasks as they arise. This rapid adaptation capability is particularly useful in applications that require frequent updates or expansions. In our experiments 4.5, we also found that using a hypernetwork can significantly enhance the generalization ability for tasks. This is because it not only retains knowledge within the domain-specific modules but also learns to generate question-targeted knowledge to be inserted into the model.
Appendix B Experimantal Settings
B.1 Background
Our task formulation follows retrieval augmented models for QA Guu et al. (2020); Sachan et al. (2021). Let $\mathcal{V}^{*}$ denote the infinite set, encompassing all potential strings over the tokens in vocabulary $\mathcal{V}$, and this includes the empty string. An instance within a QA dataset is defined as a triplet $(q,a,c)$ comprising question $q$, answer $a$, and context $c$, where $q,a,c\in\mathcal{V}^{*}$. Conventionally, the context $c$ is drawn from the knowledge corpus $\mathcal{Z}$, like Wikipedia, whereby $\mathcal{Z}\subset\mathcal{V}^{*}$.
The goal of QA is to learn a distribution function, represented as $p(a|q)$, wherein the models decode a string $a$ that serves as an abstractive answer to a given query $q$. In a closed-book setting, LLMs directly encode the given question and predict the answer Roberts et al. (2020b). Specifically, considering the context $c$ as the empty string, the reliance is solely on the model parameters, i.e., $\hat{a}=\arg\max_{a\in\mathcal{V}^{*}}p(a|q,\theta)$, where $\theta$ represents the LLMs’ parameters. However, employing a direct approach of requesting models to output answers frequently results in subpar performance, primarily attributable to omitting a substantial amount of world knowledge during the process. Therefore, a popular approach is open domain setting, which marginalizes $p(a|q,c)$ over contexts $c$ in the knowledge corpus Lewis et al. (2020); Sachan et al. (2021) or generated from models Yu et al. (2023). Given the computational infeasibility of calculating probabilities for all contexts, $p(a|q,c)$ is approximated to the sum of probabilities for top $k$ contexts, i.e., $p(a|q,c)=\sum\limits^{c_{i}\in c}\limits_{c\in TopK(q)}p(a|q,c_{i})p(c_{i}|q)$, where $Topk(q)$ denotes the set of resulting top $k$ passages after the retrieval or generated with a query $q$.
B.2 Prompts for Explicit Imagine with LLMs
The prompt for explicit imagination of the context generator to imagine a short dummy useful document is:
[…]
Table 11 shows the full prompts for zero-shot results on LLM that we use for open domain QA: NQ, TQA, WQ.
Models Docu- ments Steps Lr Batch Size
T5 0 40000 1e-4 8
LoRA-Base 0 40000 5e-4 8
IMcQA 0 50000 1e-3 8
LoRA-l 0 40000 1e-4 4
IMcQA-l 0 50000 5e-4 4
FiD-3b 0 40000 1e-4 2
LoRA-3b 0 40000 1e-4 4
IMcQA 0 50000 1e-4 1
LoRA-Base 0 40000 5e-4 8
IMcQA 0 50000 1e-3 8
LoRA-l 0 40000 1e-4 4
IMcQA-l 0 50000 5e-4 4
LoRA-3b 0 40000 1e-4 2
IMcQA-3b 0 50000 1e-4 1
IMcQA 10 50000 5e-4 1
IMcQA-l 10 50000 5e-4 1
FiD-3b 10 40000 1e-4 1
IMcQA-3b 10 50000 1e-4 1
Table 5: Hyperparameter Settings.
Models # Docu- NQ TQA WQ
ments NQ TQA WQ NQ TQA WQ NQ TQA WQ
T5 0 22.16 3.18 4.12 2.65 21.8 3.15 0.88 2.95 28.3
LoRA-Base 0 16.17 4.71 6.89 3.15 21.16 0.00 1.33 3.04 26.38
IMcQA 0 23.89 6.21 10.94 5.31 22.69 6.30 3.23 5.10 30.31
LoRA-Base 1 37.17 45.20 15.62 19.57 55.37 12.50 14.15 30.89 28.88
IMcQA 1 40.14 46.61 18.92 24.78 60.75 12.82 17.70 35.24 41.06
FiD 10 46.81 53.93 24.02 28.57 63.32 17.83 18.81 41.88 41.78
IMcQA 10 47.01 55.74 24.13 31.77 64.95 19.52 24.43 48.10 46.36
T5-l 0 28.5* 3.18 4.12 2.65 28.7* 3.15 0.88 2.95 30.6*
LoRA-l 0 17.70 7.49 8.66 3.54 23.87 4.72 0.00 5.65 29.13
IMcQA-l 0 29.32 10.17 14.06 7.02 30.11 7.81 2.65 7.06 32.68
LoRA-l 1 37.61 48.50 20.71 20.54 62.71 14.81 15.36 33.83 39.37
IMcQA-l 1 42.32 54.80 22.05 26.11 65.48 18.11 18.58 47.46 45.28
FiD-l 10 46.7* 57.93 25.12 34.29 61.9* 19.64 27.65 53.87 48.1*
IMcQA-l 10 49.92 60.03 25.79 34.35 69.67 20.28 30.19 54.94 51.52
Table 6: OOD results. The primary row in the table header delineates the dataset trained, while the underscored secondary row demonstrates the in-distribution performance. IMcQA attains optimal performance both in-distribution and OOD under diverse document configurations.
B.3 Implementations
In this section, we describe the implementation of our experiments in detail, including the baseline methods, backbone models, and hyperparameters. Our model is built based on the T5 Roberts et al. (2020a). Differing from fine-tuning all model parameters $\theta$ of the updated Pre-trained Language Model (LLM), LoRA Hu et al. (2021) freezes all pre-trained Transformer parameters and optimizes only the parameters of each LoRA adapter. We employ LoRA to train a parameter-efficient fine-tuning baseline. Drawing from this, our approach updates only the parameters of the Hypernetwork to generate the weights for each LoRA adapter. This method is adopted based on LongLoRA’s Chen et al. (2023) recommendations and experimental findings, demonstrating improved performance when the normalization and FFN layers components are updated. This is because: 1) dynamically generating LoRA weights enhances generalization and parameter sharing, and 2) LoRA performs comparably to fine-tuning but mitigates the risk of catastrophic forgetting.
For the baseline, most of the hyperparameters are the default parameters of FiD Izacard and Grave (2021). For LoRA Hu et al. (2021), add the LoRA module only to the $\mathcal{QV}$ of the attention layers and also release the normalization and FFN layers.
We consider conducting experiments using three different sizes of T5, namely T5-base, T5-large, T5-3b, and Llama2-7B, Llama2-13B Touvron et al. (2023). Due to memory constraints and online distillation limitations, A100 supports processing 20 documents for T5-3b, while Llama2 does not support distillation. All experiments with T5-3b are conducted on 2 A100 GPUs, T5-large on 2 A6000 GPUs, and T5-Base on 2 RTX 3090 GPUs. However, experiments with Llama2-7b and 13b, except for IMcQA on 2 A100 GPUs, are tested on 8 RTX 3090 GPUs.
B.3.1 Hyperparameters
The detailed hyperparameter setting is as shown in Table 5. For the LoRA modules, we set the $\alpha$ 32 and the lora rank 32.
B.3.2 Baselines
DPR   Karpukhin et al. (2020) generates by searching for the most relevant documents through dense vector space representation.
FiD   Izacard and Grave (2021) retrieve relevant documents and send them separately to the Encoder, then fuse the information in the Decoder.
RFiD   Wang et al. (2023a) uses the encoder of FiD to distinguish between causal and incidental features, and guides the decoder to generate answers based on this distinction.
EAR   Chuang et al. (2023) significantly enhances the traditional sparse retrieval method BM25 by connecting query expansion models and retrievers.
FILCO   Wang et al. (2023d) identifies useful context based on lexical and information-theoretic methods.
GENREAD   Yu et al. (2023) prompt LLMs like InstructGPT Ouyang et al. (2022) to generate a large number of relevant documents and let the reader process them.
LoRA   We use LoRA Hu et al. (2021) to obtain an efficiently fine-tuned baseline and compare it with our method.
B.3.3 Evaluation
For QA datasets, we choose the exact match (EM) score Rajpurkar et al. (2016) as the evaluation metric. An answer is deemed correct if it aligns with any of the responses in the list of acceptable answers after normalization. Normalization involves transforming the text into lowercase, omitting articles, punctuation, and eliminating redundant spaces.
Dataset Train Dev Test
WebQ 3,417 361 2,032
NQ 79,168 8,757 3,610
TQA 78,785 8,837 11,313
Table 7: Open-Domain QA dataset statistics
B.4 Downstream Evaluation Datasets
We use the following three Open-Domain QA for the experiments (§ 4.1).
  • NaturalQuestions Kwiatkowski et al. (2019) contains questions corresponding to Google search queries. The open-domain version of this dataset is obtained by discarding answers with more than 5 tokens, each accompanied by a Wikipedia article containing the answer.
  • TriviaQA Joshi et al. (2017) contains questions gathered from trivia and quiz-league websites. The unfiltered version of TriviaQA is used for open-domain question answering, each question is accompanied by pages from web and Wikipedia searches that may contain the answer.
  • WebQuestions Berant et al. (2013) contains questions from web queries matched to corresponding entries in FreeBase Bollacker et al. (2008).
Table 7 presents detailed statistics of the dataset sizes, including the training, development, and test sets. We note that all our models are trained exclusively on the training data, and we did not include the development data in our training process. Therefore, the performance numbers reported in the paper for the dev and test data are independent of the training data.
Figure 6: Zero-Shot results (EM, %) of Llama2-13B on three open-domain QA datasets. The number in parentheses indicates the number of documents used.
Appendix C Full Experimental Results
C.1 Supervised Performance
As shown in Table 12, our initial observations indicate that regardless of the method implemented, supplying a certain quantity of related documents can expedite improvement and enhance performance in QA. FiD Izacard and Grave (2021) model outclasses all baseline models in performance. Notably, utilizing FiD-xl with a mere 10 documents yields performance on par with that attained through the use of FiD-l with 100 documents. Larger models not only encapsulate more knowledge but also demonstrate a superior ability to activate and apply this knowledge efficiently.
Additionally, in comparison with LoRA Hu et al. (2021) methods, IMcQA enhances EM scores by an average of +2.2%. In the closed-book setting, the LoRA method manifests a substantial decrease in performance, likely attributable to the inadequacy of learning sufficient knowledge via questions for storage in the LoRA module. On the other hand, IMcQA harnesses both explicit and implicit imaginations to exploit knowledge for improved outcomes. These results indicate that the knowledge stored in the LLMs’ parameters can still be further exploited.
Figure 7: Zero-Shot results (Best_Subspan EM, %) of Llama2-7B on three open-domain QA datasets.
C.2 OOD Results
Table 6 shows the full OOD results in QA. It can be observed that our method has the best OOD generalization ability on all three benchmarks. Although LoRA performs well on the in-distribution part, its performance is generally poor on OOD, with some even showing negative performance. This highlights the importance of the domain adaptability of the implicit imagination Hypernetwork in our method, which generates LoRA adapter weights based on input.
C.3 Zero-Shot Results
LLMs have limited capacity to utilize extensive context effectively and are prone to generating illusions and redundant content. Best_subspan EM assesses whether the answer is included in the output. Previous studies have corroborated that LLMs encapsulate a considerable volume of knowledge and exhibit robust performance in QA.
Here, we report the Best_Subspan_EM values of Llama2-7B and Llama2-13B on three QA datasets. From Figure 7 and Figure 8, it can be observed that Best_Subspan_EM significantly improves, but the EM values are relatively small. This indicates that LLMs may not effectively utilize retrieval documents and are prone to outputting a lot of irrelevant information. Therefore, there is an urgent need to explore efficient techniques that leverage external information and internal knowledge.
Figure 8: Zero-Shot results (Best_Subspan EM, %) of Llama2-13B on three open-domain QA datasets.
However, the model did exhibit a weak adherence to instructions, often failing to output the exact answer. Remarkably, Llama2-13B displayed a decline in EM with increased document length on the WQ dataset, whereas the Best_Subspan_EM value augmented. Contrarily, our method excelled in extracting key information by using text imagination during the compression phase.
Model NQ TriviaQA WebQ
LLaMA-2-7B
Zero-shot 8.6 14.5 2.6
DPR + ICL 18.3 32.5 15.6
DPR + RECITE Sun et al. (2023) 16.8 43.9 24.8
DPR + HICL Wang et al. (2024) 25.1 47.5 28.1
DPR + ImcQA (Ours) 33.7 44.5 31.9
Table 8: Zero-shot results of Llama2-7B
C.4 OOD and Ablation Experiment Results
Here, we supplement the experimental results of LoRA and IMcQA under supervised fine-tuning in closed-book settings and the ablation results of feedforward neural network (FFN) and Long Context Distillation (LCD). It can be observed that our method like LoRA, belongs to parameter-efficient fine-tuning, and because we share the Hypernetwork to generate LoRA adapter weights, we fine-tune fewer parameters.
From Table 9, it can be seen that releasing FFN can bring more performance improvement, possibly because adding LoRA in Attention cannot fully utilize enough knowledge Yao et al. (2022). With the support of LCD, performance is further improved, with an average increase in EM of +5%. This also proves the effectiveness of our proposed LCD. In comparison with IMcQA and LoRA, it becomes more evident that LoRA tends to transfer knowledge to the LoRA module, resulting in low generalization. Our method enhances knowledge activation through dynamic generation, showing significant effects not only in-distribution but also in OOD.
C.5 Error Analysis
Using LLM as a knowledge base inevitably leads to hallucinations, which is a significant area of research in LLM development. In our quality analysis, we sampled 100 generated documents. As shown in Table 10, we found that hallucinations occurred with a probability of 4%, while the occurrence of meaningless text, such as repeated values, was 6%. Consequently, the impact of hallucinations in our method is relatively minor.
Models # Docu- Trainable NQ TQA WQ
ments Params NQ TQA WQ NQ TQA WQ NQ TQA WQ
T5 0 220M 22.16 3.18 4.12 2.65 21.8 3.15 0.88 2.95 28.3
LoRA-Base 0 28.3M 5.43 3.15 4.02 0.00 9.60 0.00 0.22 1.77 20.47
         w FFN 0 141.5M 16.17 4.71 6.89 3.15 21.16 0.00 1.33 3.04 26.38
         w FFN & LCD 0 141.5M 21.37 2.82 6.89 1.99 17.94 3.74 0.00 2.82 32.50
IMcQA 0 26.1M 5.31 3.82 5.71 0.22 10.34 2.12 0.55 2.30 16.58
         w FFN 0 139.3M 21.05 4.52 6.50 3.51 19.08 3.15 2.11 3.84 28.17
         w FFN & LCD 0 141.5M 23.89 6.21 10.94 5.31 22.69 6.30 3.23 5.10 30.31
T5-l 0 770M 28.5* 3.18 4.12 2.65 28.7* 3.15 0.88 2.95 30.6*
LoRA-l 0 42.5M 4.42 6.50 7.87 3.98 10.03 3.94 1.99 6.71 18.11
         w FFN 0 445.1M 17.70 7.49 8.66 3.54 23.87 4.72 0.00 5.65 29.13
         w FFN & LCD 0 445.1M 28.32 4.52 10.94 5.31 25.71 6.12 1.75 4.52 29.92
IMcQA-l 0 34.8M 7.08 8.90 9.45 4.42 13.14 8.66 2.43 10.17 17.72
         w FFN 0 437.5M 23.01 8.33 11.02 3.51 20.08 3.15 3.51 5.65 31.50
         w FFN & LCD 0 437.5M 29.32 10.17 14.06 7.02 30.11 7.81 2.65 7.06 32.68
Table 9: OOD and ablation experiment results. * denotes the results are from the existing papers and LCD denotes Long Context Distillation.
Hallucinations Meaningless
4% 6%
Question: When is the next Deadpool movie being released? Question: Who got the first Nobel Prize in Physics?
Document: "Deadpool (film) Deadpool is a 2016 American superhero film based on the Marvel Comics character of the same name, produced by Marvel Studios and distributed by Walt Disney Studios Motion Pictures. Document: The Nobel Prize is not a prize in itself.
Correct answer: May 18, 2018 Correct answer: Wilhelm Conrad Röntgen
Table 10: Hallucinations and Meaningless Analysis.
Methods Prompt
CBQA Please write a high-quality answer for the given question using your knowledge. Only give me the answer and do not output any other words. Question: {question} Answer:
Retrieval Please write a high-quality answer for the given question using only the provided search results (some of which might be irrelevant). Only give me the answer and do not output any other words. Context: {context} Answer the question based on the given passages. Question: {question} Answer:
Imagine Please write a high-quality answer for the given question using your knowledge and the provided imagined compressed results (some of which might be irrelevant). Only give me the answer and do not output any other words. Imagined Context: {context} Answer the question based on your knowledge and the given imagined context. Question: {question} Answer:
Table 11: Prompts for different methods on Zero-Shot setting. CBQA denotes closed-book QA that just prompts the model with the question.
Models Reader Params # Documents NQ TriviaQA WebQ
# Closed-book Setting
T5 Roberts et al. (2020a) 220M 0 25.9 23.8 27.9
T5-l Roberts et al. (2020a) 770M 0 28.5 28.7 30.6
T5-xl Roberts et al. (2020a) 3b 0 28.30 33.92 34.43
LoRA-Base 220M 0 16.17 21.16 26.38
LoRA-l 770M 0 17.70 23.87 29.13
LoRA-xl 3b 0 23.15 32.16 35.24
IMcQA (Ours) 220M 0 23.89 22.69 30.31
IMcQA-l (Ours) 770M 0 29.32 30.11 32.68
IMcQA-xl (Ours) 3b 0 29.59 35.71 37.40
# Retrieval Augmented Generation
DPR Karpukhin et al. (2020) 110M 100 41.5 56.8 41.1
RAG Lewis et al. (2020) 400M 10 44.5 56.1 45.2
FiD Izacard and Grave (2021) 220M 100 48.2 65.0 46.71
FiD-l Izacard and Grave (2021) 770M 100 51.4 67.6 50.52
FiD-xl Izacard and Grave (2021) 3b 20 55.18 72.92 52.85
FiD-l Izacard and Grave (2021) 770M 10 46.7 61.9 48.1
FiD-xl Izacard and Grave (2021) 3b 10 50.1 66.3 50.8
EAR-l Chuang et al. (2023) 770M 10 39.6 60.0 -
EAR-xl Chuang et al. (2023) 3b 10 42.3 64.6 -
RFiD-l Wang et al. (2023a) 770M 10 48.3 63.4 -
RFiD-xl Wang et al. (2023a) 3b 10 50.5 67.8 -
FILCO-xl Wang et al. (2023d) 3b 1 44.7 59.0 -
IMcQA (Ours) 220M 10 47.01 64.95 46.36
IMcQA-l (Ours) 770M 10 49.92 69.67 51.52
IMcQA-xl (Ours) 3b 5 50.87 70.34 52.78
IMcQA-l (Ours) 770M 30 53.1 70.5 52.0
# Generation Augmented Generation
GENREAD-l (sampling) Yu et al. (2023) 770M 10 40.3 67.8 51.5
GENREAD-l (clustering) Yu et al. (2023) 770M 10 43.5 70.2 53.5
GENREAD-xl (sampling) Yu et al. (2023) 3b 10 42.6 69.6 52.6
GENREAD-xl (clustering) Yu et al. (2023) 3b 10 45.6 71.6 54.4
IMcQA (Ours) 220M 10 46.22 66.70 51.43
IMcQA-l (Ours) 770M 10 48.83 70.85 54.52
IMcQA-xl (Ours) 3b 5 49.23 72.18 55.39
# Imagination Augmented Generation (Ours)
LoRA-Base 220M 1 34.51 54.05 32.28
LoRA-l 770M 1 40.05 62.81 43.70
LoRA-xl 3b 1 44.15 66.92 48.23
IMcQA 220M 1 40.14 60.75 41.73
IMcQA-l 770M 1 42.32 65.48 45.28
IMcQA-xl 3b 1 46.51 68.38 50.45
Table 12: Full QA performances (%) of different methods on three datasets. The first part (closed-book setting) indicates that explicit documentation was not utilized; The latter three parts utilize explicit augmented documents. The best results are in bold. * means that those results are from existing papers, denotes that the number of documents is generated ($\ddagger$ indicates that the number of documents is reduced due to insufficient memory for distillation).